A) A project that provides annual cash flows of $14369 for eight years costs $77894 today. What is the NPV for the project if the required return is 7 percent? (Negative amount should be indicated by a minus sign. Round your answer to 2 decimal places. (e.g., 32.16))

B) A project that provides annual cash flows of $17545 for eight years costs $75634 today. What is the NPV for the project if the required return is 19 percent? (Negative amount should be indicated by a minus sign. Round your answer to 2 decimal places. (e.g., 32.16))

C) A project that provides annual cash flows of $11191 for eight years costs $76832 today. At what discount rate would you be indifferent between accepting the project and rejecting it? (Enter your answer as a percentage, omit the "%" sign in your response, and round your answer to 2 decimal places. For example, 0.12345 or 12.345% should be entered as 12.35.)
Answer #1

Present value of annuity= payment per period * [1-(1+i)^-n]/i  

i = interest rate per period

n = number of periods

a)

NPV = -cost + present value of cash flows

= -77894 + 14369 * [1-(1.07^-8)]/0.07

= 7907.59

b)

NPV = -cost + present value of cash flows

= -75634 + 17545 * [1-(1.19^-8)]/0.19

= -6254.65

c)

NPV =0

=>

-76832 + 11191 * [1-(1+IRR)^-8]/(IRR) = 0

using excel function RATE

IRR = 3.53%